# Predicting Energy Budgets in Droplet Dynamics: A Recurrent Neural Network Approach

### v1

@@ Acknowledgements
The authors would also like to thank the financial support given by the São Paulo Research Foundation (FAPESP) grants #2013/07375-0 and #2021/14953-6, the National Council for Scientific and Technological Development (CNPq), grants #305383/2019-1 and CAPES. The authors also acknowledge the Numerical Simulation Laboratory at FCT/UNESP for their support with cluster resources. C.M. Oishi would like to thank the insightful discussions on machine learning modeling held with the research team during his visit at the AI Institute in Dynamic Systems, University of Washington.
@@ Appendix A Drop shape extraction from an image
One of the limitations observed in most flow solvers is that defining interfaces with unusual formats is a difficult process. In the experimental environment, creating spherical or elliptical drops that are strictly symmetrical is a practically impossible task. However, when running numerical simulations to reproduce such experiments, it is often chosen to use these formats because they present parametric equations that easily define them.
In order to be able to define drops with shapes similar to those observed in experiments, thus allowing the testing of network architectures without introducing noise due to shape approximations, a method capable of extracting interfaces present in digital images was also developed in this work. To extract particles from an image, the steps shown on Fig. 12 are performed:
@@
change the image to grayscale or directly to black and white (recommended only for cases that feature easy-to-extract interfaces with little to no noise);
@@
for images converted to grayscale, apply a threshold. Based on a threshold value, image pixels are converted to white or black. If a pixel has a value lower than the threshold (L), it is changed to black (0). Otherwise, it is changed to white (255). For the example in Fig. 12, a threshold of $L=110$ was applied;
@@
approximate the contour shape by collecting a set of discretized points $(i,j)$. Since the points collected by the function are actually the indices of the pixels from the image’s contour, they need to be adjusted so that they fall in the ranges $[x_{min},x_{max}]$ and $[y_{min},y_{max}]$:
$x=\frac{j(x_{min}-x_{max})}{m}+x_{max}$
$y=\frac{i(y_{min}-y_{max})}{n}+y_{max}$
@@
As the method involves extracting points through the pixels of an image’s contours, the main restriction observed is that the extraction is limited by the image’s resolution. The more pixels an image has, the more accurate the extracted interface will be.
Furthermore, as the points are extracted from the lower left corner of each pixel, aliasing behavior is generated in the extracted points, as can be seen in Fig. 13(a).
This disturbance in the initialization of points can generate atypical behavior in the earlier steps of the simulation. To help solve the aliasing problem, a technique was applied directly to the solver.
The Trapezoidal Sub-grid Undulations Removal (TSUR), developed by [58], is a mass conserving smoothing method used to reduce aliasing and abrupt undulations in curves. The result of applying the method is shown in Fig. 13(b).
@@
In addition to the usefulness of TSUR in removing the initial aliasing of the interface, the technique can also reduce numerical noise that may appear during the simulation, generating more physically accurate approximations.
@@ Appendix B Static predictions for the dimensionless numbers using geometric data
According to the results in Figs. 11 and 14 and analysing the difference in metrics from Table 8, it is evident that the performance of the one-phase neural network, operating solely on geometrical data as inputs, is inferior to that of our two-stage prediction strategy presented in Section C.2. This comparison highlights the crucial role that is undertaken by the two-phase sequential neural network approach in enhancing the accuracy and robustness of dimensionless number predictions.
Metric
Two-phase
Direct
$R^{2}$
0.99823
0.95478
$RMSE$
1.39455
7.04277
$NRMSE$
0.0141
0.07122

### v2

@@ Acknowledgements
The authors would also like to thank the financial support given by the São Paulo Research Foundation (FAPESP) grants #2013/07375-0 and #2021/14953-6, the National Council for Scientific and Technological Development (CNPq), grants #307228/2023-1 and CAPES. The authors also acknowledge the Numerical Simulation and AI Laboratory at FCT/UNESP for their support with cluster resources. C.M. Oishi would like to thank the insightful discussions on machine learning modeling held with the research team during his visit at the AI Institute in Dynamic Systems, University of Washington.
Data Availability Statement
The algorithm employed to generate the dataset, along with the machine learning codes supporting the findings of this study, are available upon reasonable request from the corresponding author.
@@ Appendix A Drop shape extraction from an image
In experimental environments, working with perfectly spherical or elliptical droplets is practically an impossible task. However, when reproducing these experiments through numerical simulations, spherical droplets are often used as an approximation due to their simple definitions through parametric functions.
To avoid using spherical approximations, a method capable of extracting Front-Tracking interfaces from experimental images was developed in this work. To extract particles from an image, the steps shown in Fig. 14 are performed. In more detail, these steps are:
@@
change the image to grayscale or directly binarize to black and white (recommended only for cases that feature easily extractable interfaces with minimal noise);
@@
binarize the grayscale image through thresholding. Pixels are binarized depending on if their values are above or below a given threshold parameter. In Fig. 14, a threshold of $L=110$ was applied;
@@
approximate the contour shape by collecting a set of discretized points $(i,j)$. Since the points collected by the function are actually the indices of the pixels from the image contour, they need to be adjusted so that they fall within the ranges $[x_{min},x_{max}]$ and $[y_{min},y_{max}]$:
$x=\frac{j(x_{min}-x_{max})}{m}+x_{max},$ (32)
$y=\frac{i(y_{min}-y_{max})}{n}+y_{max},$ (33)
@@
As the method involves extracting points through the pixelated image contour, the main limitation to this approach is the image resolution. Additionally, as the points are extracted from the lower left corner of each pixel, aliasing behavior is generated in the extracted points, as can be seen in Fig. 15(a). This disturbance in the initialization of points can generate atypical behavior in the earlier steps of the simulation.
To help solve the aliasing problem, a technique was applied directly to the solver. The Trapezoidal Sub-grid Undulations Removal (TSUR), proposed in [Nonato2004], is a mass conserving smoothing method used to reduce aliasing and abrupt undulations in curves. The result of applying the method is shown in Fig. 15(b).
@@
In addition to the effectiveness of TSUR in removing the initial aliasing of the interface, the technique can also reduce numerical noise that may appear during the simulation, generating more physically accurate Front-Tracking solutions.
@@ Appendix B Static predictions for the dimensionless numbers using geometric data
According to the results in Figs. 13 and 16 and analyzing the difference in metrics from Table 9, it is evident that the performance of the one-phase neural network, operating solely on geometrical data as inputs, is inferior to that of our two-stage prediction strategy presented in Section C.2.
Metric Two-phase Direct
$R^{2}$ 0.99823 0.95478
$RMSE$ 1.39455 7.04277
$NRMSE$ 0.0141 0.07122
@@
This is likely related to stronger correlation between the dimensionless numbers and energy values, as the dimensionless numbers are used in the non-dimensionalization of energy terms. This observation is further supported by the Pearson correlation matrix shown in Fig. 17, which reveals a high correlation among these terms and a low correlation with the geometric data. This comparison highlights the crucial role that is undertaken by the two-phase sequential neural network approach in enhancing the accuracy and robustness of dimensionless number predictions.
Fig. 17: Pearson correlation matrix illustrating the pairwise correlation coefficients between each variable.
[…]